# Reinterpreting Landauer conductance, solving the quantum measurement problem, grand unification

### v1

@@ 2.1 Application to Landauer-Buttiker conductance formulas
The above mentioned idea of exclusion principle applied to local partial states will be now used to re-describe Landauer-Buttiker formalism. We will apply exclusion principle to local partial states at zero temperature as well as to finite temperatures. It will be thus clear that Fermi-Dirac statistics applies to local partial $k$ modes and that finds strong experimental support in the overwhelmingly successful Landauer-Buttiker conductance formulas and is explained in detail below. There are no states and no Fermi-Dirac distribution function in the reservoirs unlike what is presented in Fig. 1 as a standard practice. One typically draws a Fermi-Dirac distribution function in the reservoirs as shown in Fig. 1, and as can be seen in the book by S. Datta [1]. This is in fact un-necessary or may be a matter of convinience, but could also be related to a typical mindset that at very low temperatures the reservoirs become quantum mechanical. The general belief is that quantum mechanics is fundamental and classicality emerges. As a consequnce of this general belief, one may think that at zero temperature the whole universe will become quantum mechanical with the caveat or circular logic that zero temperature can never be reached. In this general belief system one draws a Fermi-Dirac distribution inside a reservoir. Mesoscopic experiments are done at extremely low but finite temperatures such that $k_{B}T$ is less than sub-band spacing in the leads and one should never consider the reservoirs to be at zero temperature. A classical reservoir can have a chemical potential and temperature which are just parameters and not related to a Fermi-Dirac distribution of electrons. Fermi-Dirac distribution can be however, considered once the electrons enter the leads and the system that are quantum mechanical. In the region between the reservoirs we have a scattering problem that is described in general in the framework of von-Neumann quantum mechanics [2] (see ch 4 therein) that can be easily adapted for the 1D case of Fig. 1. In the Landauer’s approach, the mathematics of quantum mechanics connects to currents in nature through $ne_{0}v$ where $n$ is the number (countable) of carriers which in this case is electrons, $e_{0}$ is the charge of an electron and $v$ is its velocity.
@@
Thus we have adopted the fact that inside the system, when $k_{B}T$ is less than sub-band spacing due to lateral confinement, we can apply zero temperature quantum mechanics. At zero temparature, all states in the conduction band is filled below $\mu_{2}$ and are also equilibrated which can also be seen as the principle of detailed balance as will be explained below. The conduction band is only in the lead because of the way in which the samples are fabricated from GaAs/AlGaAs interface. The reservoir has no Fermi energy but even classical reservoirs have a chemical potential and temperature. When the electron enters the lead then it gets a chemical potential which is also the Fermi energy.
@@
Finite tempreature : Now let us see what may happen at finite temperature when states above $\mu_{1}$ and $\mu_{2}$ has a finite probability to be occupied. The left lead will be in equilibrium with the left reservoir. Say there are $\frac{dn}{dE}$ number of relevant modes at any (all) point in the left lead and not all of them are occupied. Number of electrons at any point in the left lead will be given by the number of occupied modes, that is $\frac{dn}{dE}f(E,\mu_{1})$ where $f(E,\mu_{1})$ is the Fermi-Dirac distribution function. Similarly, number of electrons in right lead is $\frac{dn}{dE}f(E,\mu_{2})$.
@@
Assuming an exact LPDOS of $\frac{1}{hv}$ in the finite leads play a further role when we discuss charge neutrality in the leads as will be presented below. It assumes that there are wave-vectors +k+k and k-k to which we can apply Pauli exclusion principle separately. The $\pm k$ wave-vectors naturally come from the wavelets making the wave-packet which is the central object of study in quantum mechanics. But assuming that k-k states are occupied only if there is reflection (as will be done below) from the scatterer etc, and apply Pauli exclusion principle for the $\pm k$ modes separately without questioning whether they define quantum states is ad hoc. While developing a conductance formula it is not convinient to talk of energy states, and although one may argue that energy is conserved between the reservoirs, $\pm k$ belong to the same quantum state (energy state), which would mean that if +k+k state is occupied then k-k state cannot be occupied. In fact that the current through a system is given by the transmission probability through the system was known much before Landauer and it was thought, that happens only when we have sparsely filled band or very low transmission probability such that multiple electrons competing for a state is negligible. One may see [5] or can also see the more recent work of [6]. Surprise came with the success of Landauer approach for densely filled bands, transmission probability approaching the maximum limit etc. A lot of space has been given in S. Datta’s book to discuss this matter but a confusion can only be expressed in a confusing manner as can be seen on page 93 of [1].
@@
Towards a four probe Landauer conductance formula at zero temperature: Note that the central object in quantum mechanics that is a wave-packet can be decomposed into wavelets and that naturally give four different wavelets. They are +k+k mode moving forward in time, +k+k mode moving backward in time, k-k mode moving forward in time and k-k mode moving backward in time. In this section we are only talking about $\pm k$ modes in a wave-packet moving forward in time. A +k+k mode moving in forward time is physically indistinguishable from a k-k mode moving in backward time and to keep the arguements simple we only consider forward moving modes. If backward moving modes manifests (for example it does in the viscinity of a Fano resonance when the entire wave-packet can move backwards in time [2]) the arguements can be accordingly adjusted easily. Once again by zero temperature we only mean $k_{B}T$ is less than the subband spacing in the leads (see ch4 in [2]). For the two probe conductance, the voltage difference betwee the two classical reservoirs are considered. For a four probe conductance we need the voltage difference between the two leads and the leads are quantum in nature, thus leading to complicacies related to measurements in quantum mechanics. Theoretically we can always calculate a voltage difference between the two quantum leads based on the countability of electrons, and so we proceed to calculate the chemical potentials $\mu_{A}$ of left lead and $\mu_{B}$ of right lead, at zero temperature. For a completely opaque barrier we will get the situation $\mu_{A}=\mu_{1}$ and $\mu_{B}=\mu_{2}$. For a barrier that transmits, we will get a situation $\mu_{1}>\mu_{A}>\mu_{B}>\mu_{2}$ as shown in Fig. 1. First consider the right lead where we only have to consider the +k+k modes for which the number of available states at a point is given by the LPDOS, and for the relevant current carrying energy window the number of electrons is $\frac{dn}{dE}(\mu_{1}-\mu_{B})$. Of all these available states, a state will be occupied only if it receives an electron transmitted across the scatterer from the left reservoir. Therefore the number of occupied states in right lead, between $\mu_{B}$ and $\mu_{1}$ is
@@
All the k-k states will be unoccupied because the only way that these states can be occupied is if they recieve an electron from the right reservoir and at zero temperature this cannot happen in the energy interval $\mu_{B}-\mu_{2}$. Any transmitted electron from left or reflected electron from the right can only go to a +k+k state. Which means all the $\frac{dn}{dE}(\mu_{B}-\mu_{2})$ left moving states in this energy range will be unoccupied in the right lead. So the total number of unoccupied states in right lead will be
@@
The first term is for +k+k modes and all of them are occupied by electrons coming from left reservoir. The second term is for k-k modes that are occupied only if there is reflection whose probability is $|r(E)|^{2}$. Number of unoccupied stsates in left lead between $\mu_{2}$ and $\mu_{A}$ will be
@@
Here the transmission from $p$ to $q$ or from $q$ to $p$ is with the other probes removed or else $J^{u}_{qp}$ and $|t_{qp}|$ will not be related. Trnasmission to other leads will contribute to $J^{u}_{qp}$. We have also seen that the two probe conductance best captures the property of the system in terms of a sample resistance that is independent of measurements along with a constant additive factor that is the contact resistance. Thus for the two probe case, when we calculate net current flowing from lead $p$ to lead $q$, this contact resistance will cancel to give us
@@
which is a charecteristic current response for the system due to $\mu_{p}$ and $\mu_{q}$ alone with other probes removed. It is to be noted that $|t_{qp}|^{2}(\mu_{p}-0)$ is a notation for $|t_{qp}|^{2}\Delta\mu_{p}$ and that $|t_{qp}|^{2}(\mu_{p}-0)\neq|t_{qp}|^{2}(\mu_{p})-|t_{qp}|^{2}(0)$. Never the less the zero chemical potentials will not contribute because we expect that for $\mu_{p}=\mu_{q}\neq 0$ we should get $J_{qp}=0$ as a property of the system independent of measurements and hence, $t_{pq}=t_{qp}\neq 0$, which is also what we would get from quantum mechanics of a single particle. For $\mu_{p}=\mu_{q}$ there will be no current in the system even if no measurements are made. Thus
@@
Comparing with Eq. 14 one has to notice the following differences. In case of Eq. 14, without any loss of generality, we had assumed $(\mu_{1}-\mu_{2})$ is positive, implying current is flowing from lead 1 to lead 2, and that below $\mu_{2}$ there is no current. In the present case with many contacts, $(\mu_{p}-\mu_{q})$ can be positive as well as negative. In Eq. 29, $J_{qp}$ can be positive as well as negative and is a property of the system, that generally may not be measured when there will be many particles and many leads involved. In reality, with only two leads, below $\mu_{2}$, the current form 1 to 2 equals the current from 2 to 1 and this is expected from the principle of detailed balance, and this principle can safely be assumed in Eq. 14. And so for Eq. 29, we are considering the currents all the way down to zero incident energy, that is we are not assuming any detailed balance between two particular reservoirs. Only expectation is that multiple reservoirs will compete to ensure some kind of detailed balance that will come out of our analysis and tested experimentally. Indeed we will see that we will get new kinds of reciprocity relations when there are many leads, that are different from Onsager reciprocity relation that was derived from the principle of detailed balance in the thermodynamic limit. Onsager had shown that for resistivity, $\rho_{xx}(B)=\rho_{xx}(-B)$ and so one would expect that resistance will remain the same on reversing the magnetic field. But signature of this was not found for multi lead mesoscopic systems and the broad answer has to do with the fact that for mesoscopic systems we cannot define a resistivity, as stated in the introduction. For small differences like $(\mu_{p}-\mu_{q})$, we can proceed with linear response. That is we are not considering terms in $(\mu_{p}-0)^{2}$, etc, and higher powers. For only two leads, since the measured resistance is the sample resistance plus the contact resistance, we always get $t_{pq}(B)=t_{pq}(-B)$ as expected on physical grounds, because reversing the magnetic field is just looking at the sample from above or bottom, and the two probe measured conductancd as given by Eq. 14 always exhibit this, giving an impression that there may be a $\rho_{xx}(B)=\rho_{xx}(-B)$ but multiprobes will rule that out. In general $t_{pq}(B)\neq t_{pq}(-B)$ in presense of more leads than $p$ and $q$. For example, in presence of magnetic field $B$ and multiple leads, it is found
@@
If we make $U_{q}=U_{p}$ for all $q$ in Eq. 32, then we must get $J_{p}=0$ which gives the following sum rule (which will be explicitely shown below as we discuss three probe conductance).
@@
This justifies the rest of Eq. 34. Note that the $G$s are determined by the system. The voltages and currents are determined by the $G$s. Some of the voltages may be applied and the others are response determined by the response functions $G$s. Now we have descided for the system in Fig. 2, which is a current lead and which is a voltage lead. Lead 2 is the voltage lead which will measure the voltage of the floating point to which it is attached and $U_{2}$ will be adjusted such that $J_{2}$ becomes zero and this adjusted value will give the voltage of the floating point to which lead 2 is attached. While fixed $U_{1}$ and $U_{3}$ is applied to provide a bias that drives a current from 1 to 3, the basic property of a voltage probe is that it should not draw any current. One can choose to set $U_{3}=0$ without any loss of generality and write from Eq. 36.
@@
We need not consider $J_{1}$ because by knowing $J_{2}$ and $J_{3}$, we can know $J_{1}$ from Eq. 44. Since the voltage probe does not draw any current, $J_{2}=0$ and so Eq. aaaa gives back Eq. aaaa. From Eq. aaaa for $J_{3}=J_{1}$ we get
@@
This is always arbitary to a sign depending on convention that current entering the system is positive and current leaving the system is negative or vice versa. In fact Eq. 50 should be written as $J_{1}+J_{2}-J_{3}=0$. Or we can also express it as
@@ 2.2 Demonstration of $\rho_{lpd}$ from $G_{\alpha\beta}$s
While Eqs. 53 and 54 is measured by making $J_{2}=0$ by adjusting $\mu_{2}$, one can also earth the STM tip as shown in Fig. 3 and measure the coherent current from reservoir 1 to reservoir 2. This coherent current can demonstrate $\rho_{lpd}$ as will be shown in this subsection. This coherent current can be again measured by measuring scattering probabilities as is evident from Eq. 14. In axiomatic quantum mechanics, any local state will be disturbed by measurements in an unpredictable manner and the theory of quantum mechanics rests on the fact that there can only be a random outcome. There is no theory for this measurement related randomness in quantum mechanics, as opposed to the dynamics prior to measurements for which probabilities can be calculated. Once again Buttikker and his coworkers [8] could demonstrate a connection between measured currents and LPDOS in the semi-classical regime. The real issue is in the quantum regime and that remained puzzling as LPDOS could become negative and they could not make sense of it. We had shown that negative LPDOS is consistent with negative signal propagation time or time travel [2]. In the previous subsection we also showed that LPDOS can answer the ananswered questions related to why Landauer-Buttiker formalism works so well. In this sub-section, we will demonstrate that the relation between LPDOS and measured scattering probability, as derived by Buttikker and co-workers make perfect sense even when we get negative LPDOS. It will be thus clear that the possibility of a deterministic local state with a LPDOS also solve the quantum measurement problem as will be demonstrated in this sub-section. Just as in principle, one can solve Newtons equation for $10^{23}$ particles to derive classical statistical mechanics, one can in principle calculate LPDOS for a mesoscopic sample with $10^{23}$ leads attached, to deterministically predict every single electron scattering outcome. As to why the uncertainty principle will not affect this deterministic outcome is discussed in [2] (see Eqs. 3.56 and 3.57 therein with the discussion below it), but the intuitive answer is that local states do not appear in axiomatic quantum mechanics but the axioms do not rule out its possibility by Goedel’s incompleteness theorem. This applies to any other theorem that can be proved within the axiomatic framework of quantum mechanics, like for example the Bell’s theorems.
@@
For the purpose of this demonstration we consider the three probe (three leads attached to a mesoscopic sample) set up as described below with respect to Fig. 2, but the logic can be extended to arbitrary number of leads. The mesoscopic sample is the shaded region which is typically made of semi-conductor or metal. The leads $\gamma$ or $\alpha$ are current leads that may be made up of the same material. Lead $\beta$ is an STM tip that can attach to different points in the sample and is such that we can vary its proximity to the sample. The STM tip can be set up in many different ways but we will focus on the particular case considered by Buttiker and his co-workers [8]. It will be clear as to why they could only understand the semi-classical regime of that set up. That is the set up when the STM tip is weakly coupled to the sample at a point r through a tunnelling barrier.
@@
Let chemical potential of left reservoir connected to lead $\gamma$ be $\mu_{\gamma}>0$, chemical potential of right reservoir connected to lead $\alpha$ is $\mu_{\alpha}=0$, and the other end of the the lead $\beta$ is earthed. Essentially, $\mu_{\gamma}$ and $\mu_{\alpha}$ are fixed for all incident energies (or other parameters), but $\mu_{\beta}$ can floatingly adjust to zero. For incident energy $E$ such that $0=\mu_{\alpha}<E<\mu_{\gamma}$, only lead $\gamma$ injects a current according to our classical notions. For a mesoscopic sample, according to Landauer-Buttiker formalism, this defines a quantum mechanical scattering problem. Such a set up can be obtained in the laboratory and we want to address the coherent current flowing from $\gamma$ to $\alpha$. One can see [3] that in such a situation
@@
$\rho_{lpd}(E,\alpha,\textbf{r},\gamma)=-\frac{1}{2\pi}\;|s^{\prime}_{\alpha\gamma}|^{2}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$ (56)
@@
which are local partial density of states (LPDOS) for this system, where $U(\textbf{r})$ is the electrostatic potential at the point r,

### v2

@@ 2.1 Application to Landauer-Buttiker conductance formulas
The above mentioned idea of exclusion principle applied to local partial states will be now used to re-describe Landauer-Buttiker formalism. We will apply exclusion principle to local partial states at zero temperature as well as to finite temperatures. It will be thus clear that Fermi-Dirac statistics applies to local partial $k$ modes and that finds strong experimental support in the overwhelmingly successful Landauer-Buttiker conductance formulas and is explained in detail below. There are no states and no Fermi-Dirac distribution function in the reservoirs unlike what is presented in Fig. 1 as a standard practice. One typically draws a Fermi-Dirac distribution function in the reservoirs as shown in Fig. 1, and as can be seen in the book by S. Datta [1]. This is in fact un-necessary or may be a matter of convinience, but could also be related to a typical mindset that at very low temperatures the reservoirs become quantum mechanical. The general belief is that quantum mechanics is fundamental and classicality emerges. As a consequence of this general belief, one may think that at zero temperature the whole universe will become quantum mechanical with the caveat or circular logic that zero temperature can never be reached. In this general belief system one draws a Fermi-Dirac distribution inside a reservoir. Mesoscopic experiments are done at extremely low but finite temperatures such that $k_{B}T$ is less than sub-band spacing in the leads and one should never consider the reservoirs to be at zero temperature. A classical reservoir can have a chemical potential and temperature which are just parameters and not related to a Fermi-Dirac distribution of electrons. Fermi-Dirac distribution can be however, considered once the electrons enter the leads and the system that are quantum mechanical. In the region between the reservoirs we have a scattering problem that is described in general in the framework of von-Neumann quantum mechanics [2] (see ch 4 therein) that can be easily adapted for the 1D case of Fig. 1. In the Landauer’s approach, the mathematics of quantum mechanics connects to currents in nature through $ne_{0}v$ where $n$ is the number (countable) of carriers which in this case is electrons, $e_{0}$ is the charge of an electron and $v$ is its velocity.
@@
Thus we have adopted the fact that inside the system, when $k_{B}T$ is less than sub-band spacing due to lateral confinement, we can apply zero temperature quantum mechanics. At zero temperature, all states in the conduction band is filled below $\mu_{2}$ and are also equilibrated which can also be seen as the principle of detailed balance as will be explained below. The conduction band is only in the lead because of the way in which the samples are fabricated from GaAs/AlGaAs interface. The reservoir has no Fermi energy but even classical reservoirs have a chemical potential and temperature. When the electron enters the lead then it gets a chemical potential which is also the Fermi energy.
@@
Finite temperature : Now let us see what may happen at finite temperature when states above $\mu_{1}$ and $\mu_{2}$ has a finite probability to be occupied. The left lead will be in equilibrium with the left reservoir. Say there are $\frac{dn}{dE}$ number of relevant modes at any (all) point in the left lead and not all of them are occupied. Number of electrons at any point in the left lead will be given by the number of occupied modes, that is $\frac{dn}{dE}f(E,\mu_{1})$ where $f(E,\mu_{1})$ is the Fermi-Dirac distribution function. Similarly, number of electrons in right lead is $\frac{dn}{dE}f(E,\mu_{2})$.
@@
Assuming an exact LPDOS of $\frac{1}{hv}$ in the finite leads play a further role when we discuss charge neutrality in the leads as will be presented below. It assumes that there are wave-vectors +k+k and k-k to which we can apply Pauli exclusion principle separately. The $\pm k$ wave-vectors naturally come from the wavelets making the wave-packet which is the central object of study in quantum mechanics. But assuming that k-k states are occupied only if there is reflection (as will be done below) from the scatterer etc, and apply Pauli exclusion principle for the $\pm k$ modes separately without questioning whether they define quantum states is ad hoc. While developing a conductance formula it is not convenient to talk of energy states, and although one may argue that energy is conserved between the reservoirs, $\pm k$ belong to the same quantum state (energy state), which would mean that if +k+k state is occupied then k-k state cannot be occupied. In fact that the current through a system is given by the transmission probability through the system was known much before Landauer and it was thought, that happens only when we have sparsely filled band or very low transmission probability such that multiple electrons competing for a state is negligible. One may see [5] or can also see the more recent work of [6]. Surprise came with the success of Landauer approach for densely filled bands, transmission probability approaching the maximum limit etc. A lot of space has been given in S. Datta’s book to discuss this matter but a confusion can only be expressed in a confusing manner as can be seen on page 93 of [1].
@@
Towards a four probe Landauer conductance formula at zero temperature: Note that the central object in quantum mechanics that is a wave-packet can be decomposed into wavelets and that naturally give four different wavelets. They are +k+k mode moving forward in time, +k+k mode moving backward in time, k-k mode moving forward in time and k-k mode moving backward in time. In this section we are only talking about $\pm k$ modes in a wave-packet moving forward in time. A +k+k mode moving in forward time is physically indistinguishable from a k-k mode moving in backward time and to keep the arguments simple we only consider forward moving modes. If backward moving modes manifests (for example it does in the viscinity of a Fano resonance when the entire wave-packet can move backwards in time [2]) the arguments can be accordingly adjusted easily. Once again by zero temperature we only mean $k_{B}T$ is less than the subband spacing in the leads (see ch4 in [2]). For the two probe conductance, the voltage difference between the two classical reservoirs are considered. For a four probe conductance we need the voltage difference between the two leads and the leads are quantum in nature, thus leading to complicacies related to measurements in quantum mechanics. Theoretically we can always calculate a voltage difference between the two quantum leads based on the countability of electrons, and so we proceed to calculate the chemical potentials $\mu_{A}$ of left lead and $\mu_{B}$ of right lead, at zero temperature. For a completely opaque barrier we will get the situation $\mu_{A}=\mu_{1}$ and $\mu_{B}=\mu_{2}$. For a barrier that transmits, we will get a situation $\mu_{1}>\mu_{A}>\mu_{B}>\mu_{2}$ as shown in Fig. 1. First consider the right lead where we only have to consider the +k+k modes for which the number of available states at a point is given by the LPDOS, and for the relevant current carrying energy window the number of electrons is $\frac{dn}{dE}(\mu_{1}-\mu_{B})$. Of all these available states, a state will be occupied only if it receives an electron transmitted across the scatterer from the left reservoir. Therefore the number of occupied states in right lead, between $\mu_{B}$ and $\mu_{1}$ is
@@
All the k-k states will be unoccupied because the only way that these states can be occupied is if they receive an electron from the right reservoir and at zero temperature this cannot happen in the energy interval $\mu_{B}-\mu_{2}$. Any transmitted electron from left or reflected electron from the right can only go to a +k+k state. Which means all the $\frac{dn}{dE}(\mu_{B}-\mu_{2})$ left moving states in this energy range will be unoccupied in the right lead. So the total number of unoccupied states in right lead will be
@@
The first term is for +k+k modes and all of them are occupied by electrons coming from left reservoir. The second term is for k-k modes that are occupied only if there is reflection whose probability is $|r(E)|^{2}$. Number of unoccupied states in left lead between $\mu_{2}$ and $\mu_{A}$ will be
@@
Here the transmission from $p$ to $q$ or from $q$ to $p$ is with the other probes removed or else $J^{u}_{qp}$ and $|t_{qp}|$ will not be related. Transmission to other leads will contribute to $J^{u}_{qp}$. We have also seen that the two probe conductance best captures the property of the system in terms of a sample resistance that is independent of measurements along with a constant additive factor that is the contact resistance. Thus for the two probe case, when we calculate net current flowing from lead $p$ to lead $q$, this contact resistance will cancel to give us
@@
which is a characteristic current response for the system due to $\mu_{p}$ and $\mu_{q}$ alone with other probes removed. It is to be noted that $|t_{qp}|^{2}(\mu_{p}-0)$ is a notation for $|t_{qp}|^{2}\Delta\mu_{p}$ and that $|t_{qp}|^{2}(\mu_{p}-0)\neq|t_{qp}|^{2}(\mu_{p})-|t_{qp}|^{2}(0)$. Never the less the zero chemical potentials will not contribute because we expect that for $\mu_{p}=\mu_{q}\neq 0$ we should get $J_{qp}=0$ as a property of the system independent of measurements and hence, $t_{pq}=t_{qp}\neq 0$, which is also what we would get from quantum mechanics of a single particle. For $\mu_{p}=\mu_{q}$ there will be no current in the system even if no measurements are made. Thus
@@
Comparing with Eq. 14 one has to notice the following differences. In case of Eq. 14, without any loss of generality, we had assumed $(\mu_{1}-\mu_{2})$ is positive, implying current is flowing from lead 1 to lead 2, and that below $\mu_{2}$ there is no current. In the present case with many contacts, $(\mu_{p}-\mu_{q})$ can be positive as well as negative. In Eq. 29, $J_{qp}$ can be positive as well as negative and is a property of the system, that generally may not be measured when there will be many particles and many leads involved. In reality, with only two leads, below $\mu_{2}$, the current form 1 to 2 equals the current from 2 to 1 and this is expected from the principle of detailed balance, and this principle can safely be assumed in Eq. 14. And so for Eq. 29, we are considering the currents all the way down to zero incident energy, that is we are not assuming any detailed balance between two particular reservoirs. Only expectation is that multiple reservoirs will compete to ensure some kind of detailed balance that will come out of our analysis and tested experimentally. Indeed we will see that we will get new kinds of reciprocity relations when there are many leads, that are different from Onsager reciprocity relation that was derived from the principle of detailed balance in the thermodynamic limit. Onsager had shown that for resistivity, $\rho_{xx}(B)=\rho_{xx}(-B)$ and so one would expect that resistance will remain the same on reversing the magnetic field. But signature of this was not found for multi lead mesoscopic systems and the broad answer has to do with the fact that for mesoscopic systems we cannot define a resistivity, as stated in the introduction. For small differences like $(\mu_{p}-\mu_{q})$, we can proceed with linear response. That is we are not considering terms in $(\mu_{p}-0)^{2}$, etc, and higher powers. For only two leads, since the measured resistance is the sample resistance plus the contact resistance, we always get $t_{pq}(B)=t_{pq}(-B)$ as expected on physical grounds, because reversing the magnetic field is just looking at the sample from above or bottom, and the two probe measured current as given by Eq. 14 always exhibit this, giving an impression that there may be a $\rho_{xx}(B)=\rho_{xx}(-B)$ but multiprobes will rule that out. In general $t_{pq}(B)\neq t_{pq}(-B)$ in presense of more leads than $p$ and $q$. For example, in presence of magnetic field $B$ and multiple leads, it is found
@@
If we make $U_{q}=U_{p}$ for all $q$ in Eq. 32, then we must get $J_{p}=0$ which gives the following sum rule (which will be explicitly shown below as we discuss three probe conductance).
@@
This justifies the rest of Eq. 34. Note that the $G$s are determined by the system. The voltages and currents are determined by the $G$s. Some of the voltages may be applied and the others are response determined by the response functions $G$s. Now we have decided for the system in Fig. 2, which is a current lead and which is a voltage lead. Lead 2 is the voltage lead which will measure the voltage of the floating point to which it is attached and $U_{2}$ will be adjusted such that $J_{2}$ becomes zero and this adjusted value will give the voltage of the floating point to which lead 2 is attached. While fixed $U_{1}$ and $U_{3}$ is applied to provide a bias that drives a current from 1 to 3, the basic property of a voltage probe is that it should not draw any current. One can choose to set $U_{3}=0$ without any loss of generality and write from Eq. 36.
@@
We need not consider $J_{1}$ because by knowing $J_{2}$ and $J_{3}$, we can know $J_{1}$ from Eq. 44. Since the voltage probe does not draw any current, $J_{2}=0$ and so Eq. 51 gives back Eq. 50. From Eq. 52 for $J_{3}=J_{1}$ we get
@@
This is always arbitrary to a sign depending on convention that current entering the system is positive and current leaving the system is negative or vice versa. In fact Eq. 44 may be written as $J_{1}+J_{2}-J_{3}=0$. Or we can also express it as
@@ 2.2 Demonstration of $\rho_{lpd}$ from $G_{\alpha\beta}$s
While Eqs. 53 and 54 is measured by making $J_{2}=0$ by adjusting $\mu_{2}$, one can also earth the STM tip as shown in Fig. 3 and measure the coherent current from reservoir 1 to reservoir 2. This coherent current can demonstrate $\rho_{lpd}$ as will be shown in this subsection. This coherent current can be again measured by measuring scattering probabilities as is evident from Eq. 14. In axiomatic quantum mechanics, any local state will be disturbed by measurements in an unpredictable manner and the theory of quantum mechanics rests on the fact that there can only be a random outcome. There is no theory for this measurement related randomness in quantum mechanics, as opposed to the dynamics prior to measurements for which probabilities can be calculated. A good example to see this are what is known as snapshots of chaotic billiards. Once again Buttikker and his coworkers [8] could demonstrate a connection between measured currents and LPDOS in the semi-classical regime. The real issue is in the quantum regime and that remained puzzling as LPDOS could become negative and they could not make sense of it. We had shown that negative LPDOS is consistent with negative signal propagation time or time travel [2]. In the previous subsection we also showed that LPDOS can answer the unanswered questions related to why Landauer-Buttiker formalism works so well. In this sub-section, we will demonstrate that the relation between LPDOS and measured scattering probability, as derived by Buttikker and co-workers make perfect sense even when we get negative LPDOS. It will be thus clear that the possibility of a deterministic local state with a LPDOS also solve the quantum measurement problem as will be demonstrated in this sub-section. Just as in principle, one can solve Newtons equation for $10^{23}$ particles to derive classical statistical mechanics, one can in principle calculate LPDOS for a mesoscopic sample with $10^{23}$ leads attached, to deterministically predict every single electron scattering outcome. As to why the uncertainty principle will not affect this deterministic outcome is discussed in [2] (see Eqs. 3.56 and 3.57 therein with the discussion below it), but the intuitive answer is that local states do not appear in axiomatic quantum mechanics but the axioms do not rule out its possibility by Goedel’s incompleteness theorem. This applies to any other theorem that can be proved within the axiomatic framework of quantum mechanics, like for example the Bell’s theorems.
@@
For the purpose of this demonstration we consider the three probe (three leads attached to a mesoscopic sample) set up as described below with respect to Fig. 2, but the logic can be extended to arbitrary number of leads. The mesoscopic sample is the shaded region which is typically made of semi-conductor or metal. The leads $\gamma$ or $\alpha$ are current leads that may be made up of the same material. Lead $\beta$ is an STM tip that can attach to different points in the sample and is such that we can vary its proximity to the sample. The STM tip can be set up in many different ways but we will focus on the particular case considered by Buttiker and his co-workers [8]. It will be clear as to why they could only understand the semi-classical regime of that set up. That is the set up when the STM tip is weakly coupled to the sample at a point r through a tunneling barrier.
@@
Let chemical potential of left reservoir connected to lead $\gamma$ be $\mu_{\gamma}>0$, chemical potential of right reservoir connected to lead $\alpha$ is $\mu_{\alpha}=0$, and the other end of the the lead $\beta$ is earthed. Essentially, $\mu_{\gamma}$ and $\mu_{\alpha}$ are fixed for all incident energies (or other parameters), but $\mu_{\beta}$ can floatingly adjust to zero as the potential at the point r where the STM tip touches the sample depend on many parameters. For incident energy $E$ such that $0=\mu_{\alpha}<E<\mu_{\gamma}$, only lead $\gamma$ injects a current according to our classical notions. For a mesoscopic sample, according to Landauer-Buttiker formalism, this defines a quantum mechanical scattering problem. Such a set up can be obtained in the laboratory and we want to address the coherent current flowing from $\gamma$ to $\alpha$. One can derive [3] that in such a situation
@@
$\rho_{lpd}(E,\alpha,\textbf{r},\gamma)=-\frac{1}{2\pi e_{0}}\;|s^{\prime}_{\alpha\gamma}|^{2}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$ (56)
@@
which are local partial density of states (LPDOS) for this system, where $U(\textbf{r})$ is the electrostatic potential at the point r and $e_{0}$ is the charge of the particles.